OpenStudy (anonymous):

The price of products may increase due to inflation and decrease due to depreciation. Derek is studying the change in the price of two products, A and B, over time. The price f(x), in dollars, of product A after x years is represented by the function below. f(x) = 72(1.25)x Part A: Is the price of product A increasing or decreasing and by what percentage per year? Justify your answer. Part B: The table below shows the price f(t), in dollars, of product B after t years:

OpenStudy (anonymous):

@jcpd910 Help again please, Last one and im done lol

OpenStudy (anonymous):

I Think A is increase

OpenStudy (anonymous):

But dk %

OpenStudy (anonymous):

f(x) = 72(1.25)x Okay, we know that it is an increase because 72 * 1.25 is 90.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Wait, one second...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ah, the difference is 18, so we divide 18 by 72 and multiply by 100. The percent increase is 25%.

OpenStudy (anonymous):

So, to answer your question: The price of Product A is increasing by 25% each year. We can justify this by multiplying 72 by 1.25, and then finding the difference between 72 and the sum of our equation. We then divide the difference by 72, and multiply by 100%. The product is our percent increase.

OpenStudy (anonymous):

I don't see the table the next part is referring to.

OpenStudy (anonymous):

ok sorry back

OpenStudy (anonymous):

I need the table.

OpenStudy (anonymous):

@jcpd910 how do i upload

OpenStudy (anonymous):

And what's the question?

OpenStudy (anonymous):

t (number of years) 1 2 3 4 f(t) (price in dollars) 65 84.5 109.85 142.81

OpenStudy (anonymous):

Part B

OpenStudy (anonymous):

The table below shows the price f(t), in dollars, of product B after t years:

OpenStudy (anonymous):

Which product recorded a greater percentage change in price over the previous year? Justify your answer. @jcpd910

OpenStudy (anonymous):

This is confusing. 'Which product' is asking about both products, or comparing them, but then 'previous year' is based on only a single product. I'll do it how I think they want it...

OpenStudy (anonymous):

The price f(x), in dollars, of product A after x years is represented by the function below. f(x) = 72(1.25) Which product recorded a greater percentage change in price over the previous year? Justify your answer.

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

Correction on the function f(x) = 72(1.25)^x

OpenStudy (anonymous):

Not really, we already know it's a 25% increase each year.

OpenStudy (anonymous):

But at least now I know it's compound.

OpenStudy (anonymous):

So our next function is: f(t) = 65(x)^t

OpenStudy (anonymous):

Wait, no... better way to do this, what is it....

OpenStudy (anonymous):

Oh, durhur. 84.5 - 65 = 19.5 19.5 / 65 = 0.3 0.3 * 100% = 30%

OpenStudy (anonymous):

So Product B has a higher percent increase.

OpenStudy (anonymous):

Thank you very much person

OpenStudy (anonymous):

No problem person.

OpenStudy (anonymous):

Well cya time for some Bf4 and Cod Lol